OpenStudy (anonymous):

Solve for x be sure to show ALL STEPS Please . 10 4 5 ---------- + ------ = --------- x(x-2) x x-2

jimthompson5910 (jim_thompson5910):

The problem is \[\Large \frac{10}{x(x-2)} + \frac{4}{x} = \frac{5}{x-2}\] right?

OpenStudy (anonymous):

Yes , Correct . @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok before I attempt to solve, I should point out one thing

jimthompson5910 (jim_thompson5910):

we CANNOT divide by zero so each denominator CANNOT be zero

OpenStudy (anonymous):

Oh okay . Got it !

jimthompson5910 (jim_thompson5910):

so x cannot equal 0 and x-2 cannot equal 0 -----> x cannot equal 2

jimthompson5910 (jim_thompson5910):

now let's solve

jimthompson5910 (jim_thompson5910):

\[\Large \frac{10}{x(x-2)} + \frac{4}{x} = \frac{5}{x-2}\] \[\Large \frac{10}{x(x-2)} + \frac{4(x-2)}{x(x-2)} = \frac{5x}{x(x-2)}\] \[\Large \frac{10}{x(x-2)} + \frac{4x-8}{x(x-2)} = \frac{5x}{x(x-2)}\] \[\Large \frac{10+4x-8}{x(x-2)} = \frac{5x}{x(x-2)}\] \[\Large \frac{4x+2}{x(x-2)} = \frac{5x}{x(x-2)}\] \[\Large 4x+2 = 5x\] \[\Large 2 = 5x-4x\] \[\Large 2 = x\] \[\Large x = 2\] So the only *possible* solution is x = 2 But wait...before I said that x cannot equal 2 (since this causes a division by zero error). So this means that there are NO solutions

jimthompson5910 (jim_thompson5910):

If you didn't know about this restriction, then checking the possible solution of x = 2 would have given you 0 in the denominator somewhere, which would lead you conclude that x = 2 is not a solution at all so that check would show there are no solutions

OpenStudy (anonymous):

Thank you for breaking it down . This really helped me because i have a couple more questions like this . And Okay now i understand . Thanks a MILLION ! :)

jimthompson5910 (jim_thompson5910):

you're very welcome, glad it's starting to click

OpenStudy (skullpatrol):

Do you know why you can never divide by 0?

OpenStudy (anonymous):

Who ? Me ? @skullpatrol

OpenStudy (skullpatrol):

Yes. @Forever_Me

OpenStudy (skullpatrol):

@Forever_Me Would you like me to explain why?

OpenStudy (anonymous):

Yes please ! @skullpatrol

OpenStudy (skullpatrol):

Ok, but you have to ask questions if you don't understand something.

OpenStudy (skullpatrol):

@Forever_Me will you do that?

jimthompson5910 (jim_thompson5910):

one way to think of why you can't divide by zero if you could divide by zero, then let's say x/0 = 0 multiply both sides by 0 to get x = 0*0 which leads to x = 0 Now say you let x = 3 x/0 = 0 3/0 = 0 3 = 0 which is impossible ------------------------------------------------------- Now let's say x/0 = 1 x/0 = 1 x = 1*0 x = 0 but if x = 3 for instance, then you get yet another contradiction so these contradictions are some examples of why division by zero is undefined

OpenStudy (anonymous):

Okay @skullpatrol

OpenStudy (anonymous):

I never really understand wh we couldnt divided by zero , So thanks for explaning to me (: @jim_thompson5910

OpenStudy (anonymous):

Why*

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (skullpatrol):

Let us start with what the reciprocal of a number means: reciprocal: Two numbers whose product is 1, also called multiplicative inverses. 5 and 1/5 are reciprocals because 5*(1/5)=1

OpenStudy (skullpatrol):

What is the reciprocal of 4/5 @Forever_Me ?

OpenStudy (skullpatrol):

@Forever_Me What number times 4/5 equals 1?

OpenStudy (anonymous):

Umm would it be : 5 ?

OpenStudy (skullpatrol):

Does 5*(4/5) =?= 1

OpenStudy (skullpatrol):

Remember the definition of a reciprocal...Two numbers whose product is 1

OpenStudy (anonymous):

Does it have to be a single number (ex: 5,4,6,7 ) or can it be times by a fraction to get 1 ?

OpenStudy (skullpatrol):

It can be any number you want.

OpenStudy (skullpatrol):

So yes, it can be a fraction.

OpenStudy (skullpatrol):

It does not have to be a single "digit" like 5,4,6,7

OpenStudy (anonymous):

4/5 times 5/4 =1

OpenStudy (skullpatrol):

Correct!!! Very GOOD :-D

OpenStudy (anonymous):

:) Thankss !

OpenStudy (skullpatrol):

What is the reciprocal of 6?

OpenStudy (skullpatrol):

Remember: 5 and 1/5 are reciprocals because 5*(1/5)=1

OpenStudy (anonymous):

1/6 TIMES 6 = 1

OpenStudy (skullpatrol):

Yes! What is the reciprocal of 1000?

OpenStudy (anonymous):

1000 times 1/1000 =1 .

OpenStudy (skullpatrol):

Yes! So your answer to the question: "What is the reciprocal of 1000?" Should be 1/1000, because 1000 times 1/1000 =1

OpenStudy (skullpatrol):

@Forever_Me do you really understand this?

OpenStudy (anonymous):

Honestly i really do . You broke it down to the point were i could really understand this . (:Thank you .

OpenStudy (skullpatrol):

Ok, now comes the question: What is the reciprocal of 0?

OpenStudy (anonymous):

No solution ? Because 0=0 over 1 so the recipricol is 1 over 0 which is undefined. Right

jimthompson5910 (jim_thompson5910):

What skullpatrol is saying is this if p is the reciprocal of q, then p times q = 1 ------------------------------------------------------- but 0 times any number = 0 any number times 0 = 0 so there's no way for a*0 = 1 or 0*b = 1 to be true for any value of 'a' or 'b', which means that 0 has no reciprocal

OpenStudy (skullpatrol):

Remember this What number times 4/5 equals 1? Here the question becomes What number times 0 equals 1?

OpenStudy (skullpatrol):

Similarly here: So your answer to the question: "What is the reciprocal of 1000?" Should be 1/1000, because 1000 times 1/1000 =1 Becomes "What is the reciprocal of 0?" Should be "no-number," because 0 times "no-number" =1